### v3

# Multivariate Probabilistic Time Series Forecasting with Correlated Errors

Abstract
Accurately modeling the correlation structure of errors is essential for reliable uncertainty quantification in probabilistic time series forecasting. Recent deep learning models for multivariate time series have developed efficient parameterizations for time-varying contemporaneous covariance, but they often assume temporal independence of errors for simplicity. However, real-world data frequently exhibit significant error autocorrelation and cross-lag correlation due to factors such as missing covariates. In this paper, we present a plug-and-play method that learns the covariance structure of errors over multiple steps for autoregressive models with Gaussian-distributed errors. To achieve scalable inference and computational efficiency, we model the contemporaneous covariance using a low-rank-plus-diagonal parameterization and characterize cross-covariance through a group of independent latent temporal processes. The learned covariance matrix can be used to calibrate predictions based on observed residuals. We evaluate our method on probabilistic models built on RNN and Transformer architectures, and the results confirm the effectiveness of our approach in enhancing predictive accuracy and uncertainty quantification without significantly increasing the parameter size.
1 Introduction
Uncertainty quantification is vital in time series forecasting, especially for applications that need more detailed insights than point forecasts. Probabilistic time series forecasting with deep learning (DL) has gained attention for its ability to capture complex/nonlinear dependencies and provide the probability distribution of target variables [1]. In multivariate time series, autoregressive models are popular for probabilistic forecasting [2, 3, 4], modeling the joint one-step-ahead predictive distribution and making multistep-ahead predictions in a rolling manner. To achieve scalable learning, these models often assume that errors are independent over time. Typically, time series variables follow a Gaussian distribution $\mathbf{z}_{t}=m(\mathbf{h}_{t})+\boldsymbol{\eta}_{t}$, where $m(\cdot)$ is the mean function and $\boldsymbol{\eta}_{t}\sim\mathcal{N}(\mathbf{0},\boldsymbol{\Sigma}_{t})$ is a stochastic error process with contemporaneous covariance matrix $\boldsymbol{\Sigma}_{t}$. The assumption of time-independence implies $\operatorname{Cov}(\boldsymbol{\eta}_{s},\boldsymbol{\eta}_{t})=\mathbf{0}$, $\forall s\neq t$. This holds when the model can incorporate all correlations between successive time steps via hidden states determined by previous values. However, real-world data often violate this assumption, with residuals showing substantial cross-correlation due to omission of important covariates and model misspecification.
Modeling error autocorrelation (or cross-correlation) is a key research topic in statistical time series models. A common approach is to assume that the error series follows a dependent temporal process, such as an autoregressive integrated moving average (ARIMA) model [5]. DL models face similar challenges. Previous studies have tried to incorporate temporally-correlated errors into the training process by modifying the loss function [6, 7]. However, these methods, based on deterministic output, are not easily applicable to probabilistic forecasting models, especially in a multivariate setting. A notable innovation is the batch training method introduced by Zheng et al. [8], which trains a univariate probabilistic forecasting model using generalized least squares (GLS) loss over batched errors. This approach parameterizes a dynamic covariance matrix to characterize error autocorrelation, which is then used to calibrate the predictive distribution of time series variables. This approach consistently improves probabilistic forecasting performance compared to naive training (i.e., without considering autocorrelated errors); however, it is only applicable to univariate models, such as DeepAR [9].
In this paper, we introduce an efficient method for learning error cross-correlation in multivariate probabilistic forecasting models. We focus on deep learning models that are autoregressive with Gaussian-distributed errors. Modeling cross-correlation in multivariate models is challenging due to increased dimensionality, as the covariance matrix scales with the number of time series $N$. To address this computational challenge, we propose characterizing error cross-correlation with a group of independent latent temporal processes using a low-rank parameterization of the covariance matrix. This approach prevents the computational cost from increasing with the number of time series. Our method results in a general-purpose approach to multivariate probabilistic forecasting models, offering significantly improved predictive accuracy.
Contributions:
  1. 1.
    We introduce a plug-and-play method for training autoregressive multivariate probabilistic forecasting models with a re-designed GLS loss. (§4)
  2. 2.
    We propose an efficient parameterization of the covariance matrix of errors over multiple steps, allowing for efficient computation of its inverse and determinant using the matrix inversion and determinant lemmas. (§4.1)
  3. 3.
    The learned covariance matrix can be used to fine-tune the predictive distribution conditional on observed residuals. (§4.2)
  4. 4.
    We demonstrate that the proposed method effectively captures error cross-correlation and improves prediction quality. Importantly, these improvements are realized via a statistical formulation without significantly increasing the size of model parameters. (§5)
2 Probabilistic Time Series Forecasting
Denote $\mathbf{z}_{t}=\left[z_{1,t},\dots,z_{N,t}\right]^{\top}\in\mathbb{R}^{N}$ as the vector containing the time series variables at time step $t$, where $N$ is the number of time series. Probabilistic time series forecasting can be formulated as estimating the joint conditional distribution $p\left(\mathbf{z}_{{T+1}:{T+Q}}\mid\mathbf{z}_{{T-P+1}:{T}};\mathbf{x}_{{T-P+1}:{T+Q}}\right)$ based on the observed history $\{\mathbf{z}_{t}\}_{t=1}^{T}$, where $\mathbf{z}_{{t_{1}}:{t_{2}}}=\left[\mathbf{z}_{t_{1}},\ldots,\mathbf{z}_{t_{2}}\right]$ and $\mathbf{x}_{t}$ are known time-dependent covariates (e.g., time of day and day of week) for all future time steps. In essence, the problem can be described as predicting the time series values for $Q$ future time steps, using all available covariates and $P$ steps of historical time series values:
$p\left(\mathbf{z}_{{T+1}:{T+Q}}\mid\mathbf{z}_{{T-P+1}:{T}};\mathbf{x}_{{T-P+1}:{T+Q}}\right)=\prod\nolimits_{t=T+1}^{T+Q}p\left(\mathbf{z}_{t}\mid\mathbf{z}_{{t-P}:{t-1}};\mathbf{x}_{{t-P}:{t}}\right),$ (1)
which becomes an autoregressive model that can be used for either one-step-ahead ($Q=1$) or multistep-ahead forecasting in a rolling manner. When performing multistep-ahead forecasting, samples will be drawn in the prediction range ($t\geq T+1$) and fed back for the next time step until the end of the desired prediction range. In neural networks, the conditioning information is commonly encoded into a state vector $\mathbf{h}_{t}$. Hence, Eq. (1) can be expressed more concisely:
$p\left(\mathbf{z}_{{T+1}:{T+Q}}\mid\mathbf{z}_{{T-P+1}:{T}};\mathbf{x}_{{T-P+1}:{T+Q}}\right)=\prod\nolimits_{t=T+1}^{T+Q}p\left(\mathbf{z}_{t}\mid\mathbf{h}_{t}\right),$ (2)
where $\mathbf{h}_{t}$ is mapped to the parameters of a parametric distribution (e.g., multivariate Gaussian).
Figure 1: Contemporaneous covariance matrix $\operatorname{Cov}(\boldsymbol{\eta}_{t},\boldsymbol{\eta}_{t})$ and cross-covariance matrix $\operatorname{Cov}(\boldsymbol{\eta}_{t-\Delta},\boldsymbol{\eta}_{t}),\Delta=1,2,3$, calculated based on the one-step-ahead prediction residuals of GPVar on a batch of time series from the $\mathtt{m4\_hourly}$ dataset. For visualization clarity, covariance are clipped to the range $[0,0.6]$.
Existing autoregressive models usually assume that the error at each time step is independent and thus $\mathbf{z}_{t}$ follows a multivariate Gaussian distribution:
$\left.\mathbf{z}_{t}\mid\mathbf{h}_{t}\right.\sim\mathcal{N}\left(\boldsymbol{\mu}(\mathbf{h}_{t}),\boldsymbol{\Sigma}(\mathbf{h}_{t})\right),$ (3)
where $\boldsymbol{\mu}(\cdot)$ and $\boldsymbol{\Sigma}(\cdot)$ map $\mathbf{h}_{t}$ to the mean and covariance parameters of a multivariate Gaussian distribution. This formulation is equivalent to decomposing $\mathbf{z}_{t}=\boldsymbol{\mu}_{t}+\boldsymbol{\eta}_{t}$ with $\boldsymbol{\eta}_{t}\sim\mathcal{N}(\mathbf{0},\boldsymbol{\Sigma}_{t})$. The temporally independent error assumption corresponds to $\operatorname{Cov}(\boldsymbol{\eta}_{s},\boldsymbol{\eta}_{t})=\mathbf{0}$ for any time points $s$ and $t$ such that $s\neq t$. Fig. 1 provides an empirical example of the contemporaneous covariance matrix $\operatorname{Cov}(\boldsymbol{\eta}_{t},\boldsymbol{\eta}_{t})$ and cross-covariance matrix $\operatorname{Cov}(\boldsymbol{\eta}_{t-\Delta},\boldsymbol{\eta}_{t}),\Delta=1,2,3$. The results are calculated based on the prediction residuals of GPVar [2] on the $\mathtt{m4\_hourly}$ dataset. While multivariate models primarily focus on modeling the contemporaneous covariance, it is evident that the residuals are not independent across time, as $\operatorname{Cov}(\boldsymbol{\eta}_{t-\Delta},\boldsymbol{\eta}_{t})\neq\mathbf{0}$. The non-zero cross-covariance suggests that there still exists information left in the residuals, and we can leverage it to enhance the prediction.
3 Related Work
3.1 Probabilistic Time Series Forecasting
Probabilistic forecasting aims to model the probability distribution for target variables, unlike deterministic forecasting which produces only point estimates. There are two main approaches: parametric probability density functions (PDFs) and quantile functions [1]. For instance, MQ-RNN [10] generates quantile forecasts using a sequence-to-sequence (Seq2Seq) RNN architecture. In contrast, PDF-based approaches assume a specific distribution (e.g., Gaussian, Poisson) and use neural networks to generate the distribution parameters. DeepAR [9], for instance, uses an RNN to model hidden state transitions, while its multivariate version, GPVar [2], employs a Gaussian copula to transform observations into Gaussian variables, assuming a joint multivariate Gaussian distribution.
Neural networks can also generate probabilistic model parameters. The deep state space model (SSM) [11] uses an RNN to generate SSM parameters. The normalizing Kalman filter (NKF) [12] combines normalizing flows (NFs) with the linear Gaussian state space model (LGM) to model nonlinear dynamics and evaluate the PDF of observations. NKF uses RNNs to produce LGM parameters at each time step, then transforms the LGM output into observations using NFs. Wang et al. [13] proposed the deep factor model, which includes a deterministic global component parameterized by an RNN and a random component from any classical probabilistic model (e.g., Gaussian white noise) to represent random effects. Some methods improve expressive conditioning for probabilistic forecasting by using Transformers instead of RNNs to model latent state dynamics, thus breaking the Markovian assumption in RNNs [14]. Other approaches adopt more flexible distribution forms, including normalizing flows [3], diffusion models [4], and Copulas [15, 16]. We refer readers to Benidis et al. [1] for a recent and comprehensive review.
3.2 Modeling Correlated Errors
Error correlation in time series is a well-studied topic in the domains of econometrics and statistics [17, 5, 18]. In multivariate time series, correlation structure is characterized by contemporaneous covariance $\operatorname{Var}(\boldsymbol{\eta}_{t})=\operatorname{Cov}(\boldsymbol{\eta}_{t},\boldsymbol{\eta}_{t})$ and cross-covariance $\operatorname{Cov}(\boldsymbol{\eta}_{t-\Delta},\boldsymbol{\eta}_{t})$. Cross-covariance includes both the autocovariance of errors $\operatorname{Cov}(\eta_{i,t-\Delta},\eta_{i,t})$ and the cross-lag covariance $\operatorname{Cov}(\eta_{i,t-\Delta},\eta_{j,t})$ between all pairs of components in the multivariate series. Contemporaneous covariance captures the correlation among individual time series at a specific point in time. In the univariate setting, DeepAR [9] achieves probabilistic forecasting by modeling the contemporaneous covariance, assuming that errors are independent over time. To address autocorrelation, Sun et al. [6] re-parameterized the input and output of neural networks to model first-order error autocorrelation, effectively capturing serially correlated errors using an AR$(1)$ process. This method enhances the performance of one-step-ahead neural forecasting models, allowing joint optimization of base and error regressors, but is limited to deterministic models. In spatial modeling, Saha et al. [19] introduced the RF-GLS model, which uses random forests to estimate nonlinear covariate effects and Gaussian processes (GP) to model spatial random effects. The RF-GLS model assumes that the error process follows an AR$(p)$ process to accommodate autocorrelated errors. Zheng et al. [8] proposed training a probabilistic forecasting model with a GLS loss that explicitly models the time-varying autocorrelation of batched error terms, extending DeepAR to incorporate autocorrelated errors.
In the multivariate setting, most existing work focuses on modeling contemporaneous covariance, assuming that $\boldsymbol{\eta}_{t}$ is independently distributed, which implies $\operatorname{Cov}(\boldsymbol{\eta}_{t-\Delta},\boldsymbol{\eta}_{t})=\boldsymbol{0}$. For example, GPVar [2] generalizes DeepAR [9] to account for correlations between time series by viewing the distribution of time series variables as a Gaussian process. In Seq2Seq models, correlations can span across series and forecasting steps. Since predictions for future time steps are generated simultaneously, we refer to these correlations as contemporaneous correlations within the scope of this study. Choi et al. [20] introduced a dynamic mixture of matrix Gaussian distributions to capture contemporaneous covariance of errors in Seq2Seq models. An exception that explicitly models error cross-correlation is [7], where the authors assume that the matrix-variate error term of a multivariate Seq2Seq model follows a matrix autoregressive (AR) process with seasonal lags. However, applying this technique to probabilistic forecasting models is not straightforward.
To the best of our knowledge, our work represents the first attempt to model cross-covariance in multivariate probabilistic time series forecasting. The closest related studies are by Zheng et al. [8] and Zheng et al. [7]. Zheng et al. [8] applies GLS loss in the temporal domain to model autocorrelated errors, but their approach is tailored for univariate time series. Zheng et al. [7] models cross-covariance in multivariate forecasting models, but their method is limited to deterministic models and requires predefined seasonal lags in the error autoregressive process. Our work extends [8] to the multivariate setting, enabling the modeling of the correlation structure of multivariate errors across multiple steps. In addition, we distinguish our approach from methods that directly model the distribution of time series variables, such as Copulas [15, 16], where no decomposition of the error term is provided.
4 Our Method
Our methodology builds upon the formulation outlined in Eq. (2), employing an autoregressive model as its foundational framework. Using RNN as an example, a probabilistic forecasting model consists of two components. Firstly, it incorporates a transition model $f_{\Theta}$ to capture the dynamics of state transitions $\mathbf{h}_{t}=f_{\Theta}\left(\mathbf{h}_{t-1},\mathbf{z}_{t-1},\mathbf{x}_{t}\right)$, thus inherently having autoregressive properties. Secondly, it integrates a distribution head, represented by $\theta$, which is tasked with mapping $\mathbf{h}_{t}$ to the parameters governing the desired probability distribution. Following GPVar [2], our approach employs the multivariate Gaussian distribution as the distribution head. The time series variable can be decomposed into a deterministic mean component and a random error component $\mathbf{z}_{t}=\boldsymbol{\mu}_{t}+\boldsymbol{\eta}_{t}$, where $\boldsymbol{\eta}_{t}\sim\mathcal{N}(\mathbf{0},\boldsymbol{\Sigma}_{t})$. To efficiently model covariance $\boldsymbol{\Sigma}_{t}$ for large $N$, GPVar adopts a low-rank-plus-diagonal parameterization $\boldsymbol{\Sigma}_{t}=\boldsymbol{L}_{t}\boldsymbol{L}_{t}^{\top}+\operatorname{diag}{(\mathbf{d}_{t})}$, where $\boldsymbol{L}_{t}\in\mathbb{R}^{B\times R}$ ($R\ll B)$ and $\boldsymbol{d}_{t}\in\mathbb{R}_{+}^{B}$. Autoregressive models based on Gaussian likelihood are usually built on the assumption that $\boldsymbol{\eta}_{t}$ are independently distributed following a multivariate Gaussian distribution. The log-likelihood of the distribution serves as the loss function for optimizing the model:
$\mathcal{L}=\sum\nolimits_{t=1}^{T}\log p\left(\mathbf{z}_{t}\mid\theta\left(\mathbf{h}_{t}\right)\right)\propto\sum\nolimits_{t=1}^{T}-\frac{1}{2}[\ln\lvert\boldsymbol{\Sigma}_{t}\rvert+\boldsymbol{\eta}_{t}^{\top}\boldsymbol{\Sigma}_{t}^{-1}\boldsymbol{\eta}_{t}].$ (4)
The parameters $\theta(\mathbf{h}_{t})$ are parameterized as $(\boldsymbol{\mu}_{t},\boldsymbol{L}_{t},\mathbf{d}_{t})$, where $\boldsymbol{\mu}_{t}\in\mathbb{R}^{N}$ represents the mean vector of the distribution. $\boldsymbol{L}_{t}\in\mathbb{R}^{N\times R}$ and $\mathbf{d}_{t}\in\mathbb{R}_{+}^{N}$ correspond to the covariance factor and diagonal elements in the low-rank parameterization of the multivariate Gaussian distribution. We use shared mapping functions for all time series:
$\mu_{i}(\mathbf{h}_{i,t})$ $=\tilde{\mu}(\mathbf{h}_{i,t})=\mathbf{w}_{\mu}^{\top}\mathbf{h}_{i,t},$ (5)
$d_{i}(\mathbf{h}_{i,t})$ $=\tilde{d}(\mathbf{h}_{i,t})=\log(1+\exp(\mathbf{w}_{d}^{\top}\mathbf{h}_{i,t})),$
$\ l_{i}(\mathbf{h}_{i,t})$ $=\tilde{l}(\mathbf{h}_{i,t})=W_{l}\mathbf{h}_{i,t},$
where $\mathbf{h}_{i,t}\in\mathbb{R}^{H\times 1}$, $\mathbf{w}_{\mu}\in\mathbb{R}^{H\times 1}$, $\mathbf{w}_{d}\in\mathbb{R}^{H\times 1}$, and $W_{l}\in\mathbb{R}^{R\times H}$ are parameters. Since the parameters of mapping functions are shared for all time series, we can view $\mathbf{z}_{t}$ as a Gaussian process assessed at points $\mathbf{h}_{i,t}$. Consequently, we can leverage a random subset of time series to compute the Gaussian likelihood-based loss in each iteration. In other words, we can train the model with a substantially reduced batch size $B<N$.
4.1 Training with Correlated Errors
We expand upon the approach introduced in [8] to handle cross-correlated errors in a multivariate context by introducing time-dependent error terms $\boldsymbol{\eta}_{t}$ into the GLS loss. In many existing deep probabilistic forecasting models, such as GPVar [2], a training batch typically consists of a sample slice of $B$ time series spanning a temporal length of $P+Q$, where $P$ is the conditioning range and $Q$ is the prediction range. Gaussian likelihood is evaluated independently at each time step within the prediction range through one-step-ahead prediction. However, this direct approach overlooks the serial correlation of errors across consecutive time steps. To address this drawback, we propose a modification to the likelihood function by introducing a dynamic covariance that accommodates the temporal dependence of the error term, as illustrated in Fig. 2. To achieve this, we organize $D$ smaller slices of time series with a temporal length of $P+1$ (i.e., $Q=1$), sorted by the prediction start time in sequential order, where $D$ represents the time horizon in which we consider cross-correlation. It is evident that the new batch structure is a reconstruction of the conventional training batch, covering the same time horizon if $D=Q$. An example of the collection of target time series variables in a batch covering cross-correlation horizon $D$ is given by
$\mathbf{z}_{t-D+1}$ $=\boldsymbol{\mu}_{t-D+1}+\boldsymbol{\eta}_{t-D+1},$ (6)
$\mathbf{z}_{t-D+2}$ $=\boldsymbol{\mu}_{t-D+2}+\boldsymbol{\eta}_{t-D+2},$
$\ldots$
$\mathbf{z}_{t}$ $=\boldsymbol{\mu}_{t}+\boldsymbol{\eta}_{t},$
where for time point $t^{\prime}$, $\boldsymbol{\mu}_{t^{\prime}}$, $\boldsymbol{L}_{t^{\prime}}$ and $\mathbf{d}_{t^{\prime}}$ are the output of the model. The covariance parameterization in GPVar corresponds to
$\boldsymbol{\eta}_{t^{\prime}}=\boldsymbol{L}_{t^{\prime}}\mathbf{r}_{t^{\prime}}+\boldsymbol{\varepsilon}_{t^{\prime}},$ (7)
where $\mathbf{r}_{t^{\prime}}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{R})$ is a low-dimensional latent variable, and $\boldsymbol{\varepsilon}_{t^{\prime}}\sim\mathcal{N}(\mathbf{0},\operatorname{diag}{(\mathbf{d}_{t^{\prime}})})$ is an additional error that is independent of $\mathbf{r}_{t^{\prime}}$. We denote by $\boldsymbol{z}_{t}^{\text{bat}}=\operatorname{vec}\left(\mathbf{z}_{t-D+1:t}\right)\in\mathbb{R}^{DB}$ the collection of target time series variables in a batch, where $\operatorname{vec}(\cdot)$ is an operator that stacks all the columns of a matrix into a vector. Similarly, we have $\boldsymbol{\mu}_{t}^{\text{bat}}\in\mathbb{R}^{DB}$, $\boldsymbol{r}_{t}^{\text{bat}}\in\mathbb{R}^{DR}$, $\boldsymbol{\varepsilon}_{t}^{\text{bat}}\in\mathbb{R}^{DB}$, $\boldsymbol{d}_{t}^{\text{bat}}\in\mathbb{R}_{+}^{DB}$, and $\boldsymbol{L}_{t}^{\text{bat}}=\operatorname{blkdiag}(\{\boldsymbol{L}_{t^{\prime}}\}_{t^{\prime}=t-D+1}^{t})\in\mathbb{R}^{DB\times DR}$ has a block diagonal structure (see Fig. 2). We then have the batch-wise decomposition:
$\boldsymbol{z}_{t}^{\text{bat}}=\boldsymbol{\mu}_{t}^{\text{bat}}+\boldsymbol{L}_{t}^{\text{bat}}\boldsymbol{r}_{t}^{\text{bat}}+\boldsymbol{\varepsilon}_{t}^{\text{bat}}.$ (8)
The default GPVar model assumes the latent variable $\boldsymbol{r}_{t}$ to be temporally independent, i.e., $\operatorname{Cov}\left(\boldsymbol{r}_{s},\boldsymbol{r}_{t}\right)=\boldsymbol{0},\forall s\neq t$. However, this assumption cannot accommodate the potential cross-correlation in the error. We next introduce temporal dependencies in the latent variable in a batch by assuming $\boldsymbol{r}_{t}^{\text{bat}}\sim\mathcal{N}\left(\boldsymbol{0},\boldsymbol{C}_{t}\otimes\mathbf{I}_{R}\right)$, where $\boldsymbol{C}_{t}$ is a $D\times D$ dynamic correlation matrix. This approach corresponds to assuming that the rows in the matrix $\mathbf{r}_{t-D+1:t}=\left[\mathbf{r}_{t-D+1},\ldots,\mathbf{r}_{t}\right]$ are independent and identically distributed following $\mathcal{N}\left(\boldsymbol{0},\boldsymbol{C}_{t}\right)$. To efficiently capture dynamic patterns over time, we follow Zheng et al. [7] and express $\boldsymbol{C}_{t}$ as a dynamic weighted sum base kernel matrices: $\boldsymbol{C}_{t}=\sum_{m=1}^{M}w_{m,t}\boldsymbol{K}_{m}$, where $w_{m,t}\geq 0$ (with $\sum_{m}w_{m,t}=1$) represents the weights for each component. For simplicity, we model each component $\boldsymbol{K}_{m}$ using a kernel matrix generated from a squared-exponential (SE) kernel function, with the $(i,j)$th entry being $\boldsymbol{K}^{ij}_{m}=\exp(-\frac{(i-j)^{2}}{l_{m}^{2}})$, with different lengthscales $l$ (e.g., $l=1,2,3,\dots$). In addition, we incorporate an identity matrix into the additive structure to account for the independent noise process. This parameterization ensures that $\boldsymbol{C}_{t}$ is a positive definite symmetric matrix with unit diagonals, making it a valid correlation matrix. The weights for these components are derived from the hidden state $\mathbf{h}_{t}$ at each time step through a small neural network, with the number of nodes in the output layer set to $M$ (i.e., the number of components). A softmax layer is then used to ensure that these weights are summed up to 1. Note that the parameters of this network will be learned simultaneously with those of the base model.
Figure 2: Graphic illustration of Eq. (8), where $B$ is the number of time series in a batch, $R$ is the rank of the covariance factor, $D$ is the time window we consider cross-correlation, $P$ and $Q$ are the conditioning range and prediction range. Cross-correlation is modeled by introducing correlation in each row of matrix $\mathbf{r}_{t-D+1:t}$.
Marginalizing out $\boldsymbol{r}_{t}^{\text{bat}}$ in Eq. (8), we have $\boldsymbol{z}_{t}^{\text{bat}}\sim\mathcal{N}\left(\boldsymbol{\mu}_{t}^{\text{bat}},\boldsymbol{\Sigma}_{t}^{\text{bat}}\right)$ with covariance
$\boldsymbol{\Sigma}_{t}^{\text{bat}}=(\boldsymbol{L}_{t}^{\text{bat}})(\boldsymbol{C}_{t}\otimes\mathbf{I}_{R})(\boldsymbol{L}_{t}^{\text{bat}})^{\top}+\operatorname{diag}(\boldsymbol{d}_{t}^{\text{bat}}).$ (9)
It is not difficult to derive that for any $i,j\in\{0,1,\ldots,D-1\}$ and $i\neq j$, the proposed model can create cross-covariance $\operatorname{Cov}\left(\boldsymbol{\eta}_{t-i},\boldsymbol{\eta}_{t-j}\right)=\boldsymbol{C}_{t}^{ij}\boldsymbol{L}_{t-i}\boldsymbol{L}_{t-j}^{\top}$ for times $t-i$ and $t-j$, which is no longer $\boldsymbol{0}$. In this sense, our model can better characterize the empirical cross-covariance matrices of the residuals (see empirical examples in Fig. 1). As $\boldsymbol{\mu}_{t}^{\text{bat}}$, $\boldsymbol{L}_{t}^{\text{bat}}$, and $\boldsymbol{d}_{t}^{\text{bat}}$ are default outputs of the base probabilistic model, we can compute the overall likelihood (with overlapped data) by
$\mathcal{L}=\sum\nolimits_{t=D}^{T}\log p\left(\boldsymbol{z}_{t}^{\text{bat}}\mid\boldsymbol{\mu}_{t}^{\text{bat}},\boldsymbol{\Sigma}_{t}^{\text{bat}}\right).$ (10)
Here, computing the log-likelihood involves evaluating the inverse and the determinant of $\boldsymbol{\Sigma}_{t}^{\text{bat}}$ with size $DB\times DB$, for which a naïve implementation has a prohibitive time complexity of $\mathcal{O}\left(D^{3}B^{3}\right)$. However, it should be noted that the covariance $\boldsymbol{\Sigma}_{t}^{\text{bat}}$ in our parameterization is of the form $\boldsymbol{E}+\boldsymbol{A}\boldsymbol{C}\boldsymbol{A}^{\top}$, where $\boldsymbol{E}=\operatorname{diag}(\boldsymbol{d}_{t}^{\text{bat}})$ is diagonal, $\boldsymbol{A}=\boldsymbol{L}_{t}^{\text{bat}}$, and $\boldsymbol{C}=\boldsymbol{C}_{t}\otimes\mathbf{I}_{R}$. This allows us to leverage the Sherman–Morrison–Woodbury identity (matrix inversion lemma) and the companion matrix determinant lemma to simplify the computation:
$(\boldsymbol{E}+\boldsymbol{A}\boldsymbol{C}\boldsymbol{A}^{\top})^{-1}$ $=\boldsymbol{E}^{-1}-\boldsymbol{E}^{-1}\boldsymbol{A}(\boldsymbol{C}^{-1}+\boldsymbol{A}^{\top}\boldsymbol{E}^{-1}\boldsymbol{A})^{-1}\boldsymbol{A}^{\top}\boldsymbol{E}^{-1},$ (11)
$\det{(\boldsymbol{E}+\boldsymbol{A}\boldsymbol{C}\boldsymbol{A}^{\top})}$ $=\det{(\boldsymbol{C}^{-1}+\boldsymbol{A}^{\top}\boldsymbol{E}^{-1}\boldsymbol{A})}\det{(\boldsymbol{C})}\det{(\boldsymbol{E})}.$
Then, the likelihood calculation only requires computing the inverse and determinant of a $DR\times DR$ matrix, specifically $\boldsymbol{C}^{-1}+\boldsymbol{A}^{\top}\boldsymbol{E}^{-1}\boldsymbol{A}$. Both of these computations can be efficiently performed using its Cholesky factorization. Detailed computations are provided in Appendix §A.2.
Modeling the latent process $\boldsymbol{r}_{t}$ offers several advantages. Firstly, because $\boldsymbol{r}_{t}$ has a much lower dimension than $\boldsymbol{\varepsilon}_{t}$, modeling the cross-correlation of $\boldsymbol{r}_{t}$ results in a significantly smaller $DR\times DR$ covariance matrix compared to the $DB\times DB$ covariance matrix of $\boldsymbol{\varepsilon}_{t}$. Secondly, since $\boldsymbol{r}_{t}$ follows an isotropic Gaussian distribution, the covariance of $\boldsymbol{r}_{t}^{\text{bat}}$ can be parameterized with a Kronecker structure $\boldsymbol{C}_{t}\otimes\mathbf{I}_{R}$. This greatly simplifies the task into learning a $D\times D$ correlation matrix shared by all time series in a batch. Lastly, similar to GPVar, we can still train the model using a subset of time series in each iteration, maintaining computational efficiency.
4.2 Multistep-ahead Rolling Prediction
Autoregressive models conduct multistep-ahead forecasting in an iterative manner—the model generates a sample at each time step during prediction, serving as input for the subsequent step, and the process continues until the desired prediction range is reached. Our approach enhances this process in a similar way as in Zheng et al. [8] by offering additional calibration based on the learned correlation matrix $\boldsymbol{C}_{t}$. Assuming observations are available up to time step $t$, the conditional distribution of $\boldsymbol{\eta}_{t+1}$ given errors in the past $(D-1)$ steps, can be derived as follows:
$\boldsymbol{\eta}_{t+1}\mid\boldsymbol{\eta}_{t},\boldsymbol{\eta}_{t-1},\ldots,\boldsymbol{\eta}_{t-D+2}\sim\mathcal{N}\left(\boldsymbol{\Sigma}_{*}\boldsymbol{\Sigma}_{\text{obs}}^{-1}\boldsymbol{\eta}_{\text{obs}},\boldsymbol{\Sigma}_{t+1}-\boldsymbol{\Sigma}_{*}\boldsymbol{\Sigma}_{\text{obs}}^{-1}\boldsymbol{\Sigma}_{*}^{\top}\right),$ (12)
where $\boldsymbol{\eta}_{\text{obs}}=\operatorname{vec}\left(\left[\boldsymbol{\eta}_{t-D+2},\ldots,\boldsymbol{\eta}_{t-1},\boldsymbol{\eta}_{t}\right]\right)\in\mathbb{R}^{(D-1)B}$ represents the set of residuals, accessible at forecasting step $t+1$. Here, $\boldsymbol{\Sigma}_{\text{obs}}$ is a $(D-1)B\times(D-1)B$ partition of $\boldsymbol{\Sigma}_{t+1}^{\text{bat}}$ that gives the covariance of $\boldsymbol{\eta}_{\text{obs}}$, and $\boldsymbol{\Sigma}_{*}$ is a $B\times(D-1)B$ partition of $\boldsymbol{\Sigma}_{t+1}^{\text{bat}}$ that gives the covariance of $\boldsymbol{\eta}_{t+1}$ and $\boldsymbol{\eta}_{\text{obs}}$, i.e., $\boldsymbol{\Sigma}_{t+1}^{\text{bat}}=\begin{bmatrix}\boldsymbol{\Sigma}_{\text{obs}}&\boldsymbol{\Sigma}_{*}^{\top}\\ \boldsymbol{\Sigma}_{*}&\boldsymbol{\Sigma}_{t+1}\end{bmatrix}$. For conciseness, we omit the time index $t$ in $\boldsymbol{\Sigma}_{\text{obs}}$, $\boldsymbol{\Sigma}_{*}$ and $\boldsymbol{\eta}_{\text{obs}}$. Given that $\boldsymbol{\mu}_{t+1}$ is a deterministic output from the base model, a sample of the target variables $\tilde{\mathbf{z}}_{t+1}$ can be derived by initially drawing a sample $\tilde{\boldsymbol{\eta}}_{t+1}$ from Eq. (12) and then combining it with the predicted mean vector $\boldsymbol{\mu}_{t+1}$ by $\tilde{\mathbf{z}}_{t+1}=\boldsymbol{\mu}_{t+1}+\tilde{\boldsymbol{\eta}}_{t+1}$.
Taking the sample $\tilde{\boldsymbol{\eta}}_{t+1}$ as an observed residual, we can iteratively execute the process described in Eq. (12) to derive a trajectory of $\{\tilde{\mathbf{z}}_{t+q}\}_{q=1}^{Q}$. By repeating this procedure, we can generate a certain number of samples to characterize the predictive distribution at each time step.
5 Experiments
5.1 Evaluation of Predictive Performance
Datasets. We use widely recognized time series benchmarking datasets from GluonTS [21]. The prediction range ($Q$) for each dataset was based on their respective configurations within GluonTS. We performed a sequential split into training, validation, and testing sets for each dataset. Each dataset was standardized using the mean and standard deviation from the training set. Predictions were rescaled to their original values for computing evaluation metrics. Further details on the dataset can be found in Appendix §A.1.
Base probabilistic models. We integrate the proposed method into two distinct autoregressive models: the RNN-based GPVar [2] and the decoder-only Transformer [22]. These models are trained to generate distribution parameters as described in §4. Our approach can be applied to other autoregressive multivariate models with minimal adjustments, provided the final prediction follows a multivariate Gaussian distribution. The implementation is carried out using PyTorch Forecasting [23]. We use lagged time series values and additional features/covariates as input for both models. Details on model training (§A.3), hyperparameter tuning (§A.5), and the base model (§A.6) are provided in Appendix §A.
Dynamic correlation matrix. We introduce a limited number of additional parameters to project the state vector $\mathbf{h}_{t}$ into component weights $w_{m,t}$, which are used to generate the dynamic correlation matrix $\boldsymbol{C}_{t}$. The number of base kernels ($M$) for generating $\boldsymbol{C}_{t}$ and the associated lengthscale set $\{l_{m}\}_{m=1}^{M-1}$ are hyperparameters. We conduct a grid search over $M=2,3,4$ and two sets of lengthscales—$\{0.5,1.5,\dots\}$ and $\{1.0,2.0,\dots\}$. Models with the best validation loss are selected. These distinct lengthscales capture varying correlation decay rates, allowing the model to account for different temporal patterns. The time-varying component weights enable dynamic adaptation to evolving correlation structures over time.
Baselines. We assess the proposed method by comparing it with a baseline model trained without considering error cross-correlation (Eq. (4)). The baseline model is effectively a special case of our model with $\boldsymbol{C}_{t}=\boldsymbol{I}_{D}$. To ensure a straightforward and fair comparison, we align the cross-correlation range ($D$) with the prediction range ($Q$), ensuring identical data sampling processes for both methods. Additionally, we set $P=Q$ following the default configuration in GluonTS. We also provide VAR and GARCH as two naive baseline models (see Appendix §A.4).
Metrics. Continuous Ranked Probability Score (CRPS) [24] is used as the main evaluation metric:
$\operatorname{CRPS}(F,z)=\mathbb{E}_{F}|Z-z|-\frac{1}{2}\mathbb{E}_{F}\left|Z-Z^{\prime}\right|,$ (13)
where $F$ is the cumulative distribution function (CDF) of the predicted variable, $z$ is the observation, $Z$ and $Z^{\prime}$ are independent copies of a set of prediction samples associated with the distribution $F$. To evaluate multivariate dependencies in the time series data, we compute $\operatorname{CRPS}_{\text{sum}}$ by first summing both the forecast and ground-truth values across all time series and then calculating the $\operatorname{CRPS}$ over the resulting sums [2, 15, 16]. Other metrics are also available in Appendix §B.1.
Training dynamics. Our approach incurs additional training costs per optimization step due to the more complex likelihood function. As shown in Appendix §B.3, the training time per epoch for models using our method is generally longer than for baseline methods. However, there is no evidence that our method slows down convergence. On the contrary, it accelerates convergence for 5 out of 9 datasets for GPVar.
Benchmark results. The $\operatorname{CRPS}_{\text{sum}}$ results are presented in Table 1. Our method offers an average improvement of 13.79% for GPVar and 6.91% for Transformer. It is important to note that the degree of performance enhancement varies across different base models and datasets, influenced by factors such as the inherent data characteristics and the performance of different model architectures. The alignment between the actual correlation structure and our kernel assumption also plays a pivotal role in our method. Our method also demonstrates consistent improvement across five different metrics, with significant gains in multivariate metrics such as the energy score (Appendix §B.2). Additionally, to provide more insights, we compare the residual autocorrelation and cross-lag correlation with and without applying our method in Appendix §B.4.1, showing that our method effectively reduces cross-correlations in many scenarios. Furthermore, in Appendix §B.4.2, we separate the accuracy improvement over forecast steps for each dataset. The performance improvement is shown to be related to both the absolute time across the temporal span of the dataset (especially for time series with strong periodic patterns) and the relative time over the prediction horizon.
Table 1: $\operatorname{CRPS}_{\text{sum}}$ accuracy comparison. “w/o” denotes methods without time-dependent errors, while “w/” indicates our method. Bold values show models with time-dependent errors performing better. Mean and standard deviation are obtained from 10 runs of each model.
VAR GARCH GPVar Transformer
w/o w/ w/o w/
$\mathtt{exchange\_rate}$ 0.0033±0.0000 0.0435±0.0001 0.0068±0.0004 0.0117±0.0004 0.0055±0.0002 0.0042±0.0002
$\mathtt{solar}$ 0.7663±0.0050 0.8752±0.0015 0.7103±0.0065 0.6929±0.0039 0.4960±0.0034 0.4132±0.0027
$\mathtt{electricity}$ 0.1264±0.0006 0.2847±0.0015 0.0430±0.0005 0.0403±0.0004 0.0494±0.0004 0.0638±0.0003
$\mathtt{traffic}$ 3.5241±0.0084 0.4459±0.0005 0.1095±0.0002 0.0649±0.0002 0.0717±0.0002 0.0981±0.0002
$\mathtt{wikipedia}$ 26.2025±0.0389 0.6699±0.0045 0.1745±0.0008 0.0743±0.0009 0.0841±0.0013 0.0500±0.0005
$\mathtt{m4\_hourly}$ 0.2352±0.0008 0.2758±0.0006 0.0613±0.0004 0.0358±0.0002 0.0651±0.0004 0.0616±0.0003
$\mathtt{m1\_quarterly}$ N/A N/A 0.3942±0.0030 0.3538±0.0017 0.4448±0.0027 0.4367±0.0028
$\mathtt{pems03}$ 0.0598±0.0002 0.3202±0.0007 0.0503±0.0001 0.0491±0.0002 0.0490±0.0001 0.0386±0.0001
$\mathtt{uber\_hourly}$ N/A N/A 0.0342±0.0006 0.0222±0.0004 0.0632±0.0003 0.0513±0.0005
avg. rel. impr. 13.79% avg. rel. impr. 6.91%
Figure 3: (a) Component weights for generating $\boldsymbol{C}_{t}$ for a batch of time series ($B=8$) from the $\mathtt{m4\_hourly}$ dataset obtained by the GPVar model. Parameters $w_{0},w_{1},w_{2}$ represent the component weights of the kernel matrices associated with lengthscales $l=0.5,1.5,2.5$, and $w_{3}$ is the component weight of the identity matrix. Shaded areas distinguish different days; (b) The autocorrelation function (ACF) indicated by the correlation matrix $\boldsymbol{C}_{t}$ at 17:00. Given the rapid decay of the ACF, we only plot 12 lags to enhance visualization; (c) The corresponding covariance matrix of the associated target variables $\boldsymbol{\Sigma}_{t}^{\text{bat}}$ at 17:00. A zoom-in view of a $3B\times 3B$ region is illustrated in the plot, where the diagonal blocks represent $B\times B$ covariance matrices $\boldsymbol{\Sigma}_{t^{\prime}}$ of $\mathbf{z}_{t^{\prime}}$ over three consecutive time steps. The off-diagonal blocks describe the cross-covariance $\operatorname{Cov}(\mathbf{z}_{t-\Delta},\mathbf{z}_{t})$, $\forall\Delta\neq 0$. For visualization clarity, covariance values are clipped to the range $[0,0.03]$.
5.2 Model Interpretation
Our method captures error cross-correlation through the dynamic construction of a covariance matrix. This is achieved by combining kernel matrices with varying lengthscales in a dynamically weighted sum. A small lengthscale corresponds to short-range positive correlations, while a large lengthscale captures positive correlations over longer lags.
In Fig. 3, we depict the dynamic component weights and the resulting autocorrelation function (the first row of the correlation matrix $\boldsymbol{C}_{t}$) for a batch of time series from the $\mathtt{m4\_hourly}$ dataset spanning a four-day window. We also provide the covariance matrix of $\boldsymbol{z}_{t}^{\text{bat}}$ using the correlation matrix and model outputs at a specific time of day. We observe that the component weight $w_{3}$, corresponding to the identity matrix, dominates throughout the entire observation period. This suggests that the error correlation is generally mild over time. This behavior is influenced by the Kronecker structure used to parameterize the covariance over the low-dimensional latent variables $\boldsymbol{r}_{t}$, which assumes that all latent processes share the same autocorrelation structure. Given the Kronecker structure, the model tends to learn the mildest temporal correlation among the time series in a batch.
Moreover, we observe that the dynamic component weights adjust the correlation strengths. Specifically, when the weight assigned to the identity matrix ($w_{3}$) increases, the error process tends to be more independent. In contrast, when the weights assigned to the other kernel matrices ($w_{0}$, $w_{1}$, and $w_{2}$) are larger, the error process becomes more correlated, as the kernel matrices with different lengthscales combine to formulate a specific correlation structure. Fig. 3(a) demonstrates pronounced daily patterns in temporal correlation, particularly when errors exhibit increased correlation around 17:00 each day. The corresponding autocorrelation function is shown in Fig. 3(b). Fig. 3(c) illustrates the corresponding covariance matrix of the associated target variables within the cross-correlation horizon. The diagonal blocks represent the contemporaneous covariance $\boldsymbol{\Sigma}_{t}$ of $\mathbf{z}_{t}$ at each time step, while the off-diagonal blocks represent the cross-covariance $\operatorname{Cov}(\mathbf{z}_{t-\Delta},\mathbf{z}_{t})$ for $\forall\Delta\neq 0$, which can be captured by our approach. The zoomed-in view provides a $3B\times 3B$ region that illustrates the cross-covariance within two lags. We observe that the cross-covariance is most pronounced at lag 1, consistent with the observation in Fig. 3(a) that the component weight $w_{0}$, assigned to the base kernel matrix with lengthscale $l=0.5$, is more pronounced than $w_{1}$ and $w_{2}$.
6 Conclusion and Broader Impacts
This paper presents a novel approach to addressing error cross-correlation in multivariate probabilistic time series forecasting, specifically for models with autoregressive properties and Gaussian distribution heads. We construct a dynamic covariance matrix using a small set of independent and identically distributed latent temporal processes. These latent processes effectively model temporal correlation and integrate seamlessly into the base model, where the contemporaneous covariance is parameterized by a low-rank-plus-diagonal structure. This approach enables the modeling and prediction of a time-varying covariance matrix for the target time series variable. The experiment results demonstrate its effectiveness in enhancing uncertainty quantification.
Our contributions are two-fold. First, our approach relaxes the time-independent error assumption during the training process for probabilistic forecasting models, addressing the reality that residuals are typically not time-independent. Second, the learned cross-correlation enhances multistep-ahead predictions by refining the distribution output at each forecasting step. These enhancements to existing models have broader implications for fields such as finance, healthcare, and energy, where improved forecasts and uncertainty quantification can lead to better risk management, resource allocation, and operational efficiency.
There are several avenues for future research. Firstly, the Kronecker structure $\boldsymbol{C}_{t}\otimes\mathbf{I}_{R}$ for the covariance matrix of the latent variable $\boldsymbol{r}_{t}^{\text{bat}}$ may be too restrictive for multivariate time series problems. This structure assumes that the rows (i.e., latent temporal processes) in the matrix $\left[\mathbf{r}_{t-D+1},\ldots,\mathbf{r}_{t}\right]$ are identically distributed following $\mathcal{N}\left(\boldsymbol{0},\boldsymbol{C}_{t}\right)$. Exploring more flexible covariance structures, such as employing different $\boldsymbol{C}_{r,t}$ matrices for each of the latent temporal processes, could be a promising direction for further investigation. Secondly, the parameterization of $\boldsymbol{C}_{t}$ could be expanded. Instead of using SE kernels, $\boldsymbol{C}_{t}$ could be parameterized as fully learnable positive definite symmetric Toeplitz matrices. For example, an AR($p$) process has a covariance structure with a Toeplitz form, allowing for the modeling of negative correlations. This alternative approach could offer greater flexibility in capturing complex correlation patterns in multivariate time series data.
Acknowledgments and Disclosure of Funding
We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC). V. Z. Zheng acknowledges the support received from the FRQNT B2X Doctoral Scholarship Program.
Appendix
A Experimental Details A
A.1 Datasets A.1
A.2 Multivariate Likelihood with Autocorrelated Errors A.2
A.3 Training Procedure A.3
A.4 Naive Baseline Description A.4
A.5 Hyperparameter Search A.5
A.6 Base Model Description and Input Features A.6
B Metrics and Additional Results B
B.1 Metric Definition B.1
B.1.1 Continuous Ranked Probability Score B.1.1
B.1.2 Quantile Loss B.1.2
B.1.3 Energy Score B.1.3
B.1.4 Root Relative Mean Squared Error B.1.4
B.2 Results on Other Metrics B.2
B.3 Training Dynamics B.3
B.4 Additional Model Interpretation B.4
B.4.1 Comparison of Residual Correlation B.4.1
B.4.2 Performance Breakdown at Each Forecast Step B.4.2
B.5 Alternative Parametrization of $\boldsymbol{C}_{t}$ B.5
B.5.1 Learnable Lengthscales B.5.1
B.5.2 Using Autocorrelations of an AR$(p)$ process B.5.2
B.6 Qualitative Results on Forecasting B.6
Appendix A Experimental Details
A.1 Datasets
We performed experiments on a diverse set of real-world datasets obtained from GluonTS [21]. These datasets include:
  • $\mathtt{electricity}$ [25]: hourly electricity consumption data collected from a total of 370 households over the period spanning from January 2012 to June 2014.
  • $\mathtt{m4\_hourly}$ [26]: Hourly time series data from various domains, covering microeconomics, macroeconomics, finance, industry, demographics, and various other fields, is sourced from the M4-competition.
  • $\mathtt{exchange\_rate}$ [27]: Daily exchange rate information for eight different countries spanning the period from 1990 to 2016.
  • $\mathtt{m1\_quarterly}$ [28]: Quarterly time series data spanning seven different domains.
  • $\mathtt{pems03}$ [29]: Traffic flow records obtained from Caltrans District 3 and accessed through the Caltrans Performance Measurement System (PeMS). The records are aggregated at a 5-minute interval.
  • $\mathtt{solar}$ [27]: Hourly time series representing solar power production data in the state of Alabama for the year 2006.
  • $\mathtt{traffic}$ [30]: Hourly traffic occupancy rates recorded by sensors installed in the San Francisco freeway system between January 2008 and June 2008.
  • $\mathtt{uber\_hourly}$ [31]: Hourly time series of Uber pickups in New York City spanning from February to July 2015.
  • $\mathtt{wikipedia}$ [32]: Daily page views for 2,000 Wikipedia pages spanning from January 2012 to March 2014.
These datasets are widely employed for benchmarking time series forecasting models, following their default configurations in GluonTS, including granularity, prediction range ($Q$) and the number of rolling evaluations. For each dataset, we performed a sequential split into training, validation, and testing sets, with the temporal length of the validation set matching that of the testing set. The temporal length of the testing set was determined by considering the prediction range and the required number of rolling evaluations. For instance, the testing horizon for the $\mathtt{traffic}$ dataset is computed as $24+7-1=30$ time steps. As a result, the model will predict 24 steps ($Q$) sequentially, with 7 distinct consecutive prediction start timestamps, also known as 7 forecast instances. In our experiments, we align the conditioning range ($P$) with the prediction range ($Q$), maintaining consistency with the default setting in GluonTS. For simplicity, we set the autocorrelation horizon ($D$) to also match the prediction range ($Q$). Essentially, in this paper, we have $P=Q=D$. Each dataset was standardized using the mean and standard deviation from the training set. Predictions were rescaled to their original values for computing evaluation metrics. The statistics of all datasets are summarized in Table 2.
Table 2: Dataset summary.
Dataset Granularity # of time series # of time steps $Q$ Rolling evaluation
$\mathtt{electricity}$ hourly 370 5,857 24 7
$\mathtt{m4\_hourly}$ hourly 414 1,008 48 7
$\mathtt{exchange\_rate}$ workday 8 6,101 30 5
$\mathtt{m1\_quarterly}$ quarterly 281 48 8 1
$\mathtt{pems03}$ 5min 358 26,208 12 24
$\mathtt{solar}$ hourly 137 7,033 24 7
$\mathtt{traffic}$ hourly 963 4,025 24 7
$\mathtt{uber\_hourly}$ hourly 262 8,343 24 7
$\mathtt{wikipedia}$ daily 2,000 792 30 5
A.2 Multivariate Likelihood with Correlated Errors
The probability density function of a multivariate normal distribution considering autocorrelated errors described in §4.1 is defined in Eq. (14). For simplicity, we omit the subscript $t$ and superscript bat for all notations:
$f(\boldsymbol{z})=(2\pi)^{-B/2}\lvert\boldsymbol{\Sigma}\rvert^{-1/2}\exp{\left(-\frac{1}{2}\left(\boldsymbol{z}-\boldsymbol{\mu}\right)^{\top}\boldsymbol{\Sigma}^{-1}\left(\boldsymbol{z}-\boldsymbol{\mu}\right)\right)}.$ (14)
We use the negative log likelihood (NLL) of an observed $\boldsymbol{z}$ as the loss function for training our model. The NLL can be calculated as the negative log of the probability density function in Eq. (14):
$\mathcal{L}_{NLL}=-\ln{L\left(\boldsymbol{z}\right)}=\frac{1}{2}\left[\ln\lvert\boldsymbol{\Sigma}\rvert+\left(\boldsymbol{z}-\boldsymbol{\mu}\right)^{\top}\boldsymbol{\Sigma}^{-1}\left(\boldsymbol{z}-\boldsymbol{\mu}\right)+B\ln\left(2\pi\right)\right],$ (15)
where $B$ is the number of time series in a batch, and the covariance matrix is parameterized as $\boldsymbol{\Sigma}=\boldsymbol{L}(\boldsymbol{C}\otimes\mathbf{I}_{R})\boldsymbol{L}^{\top}+\boldsymbol{E}$, $\boldsymbol{L}\in\mathbb{R}^{DB\times DR}$ the covariance factor, $\boldsymbol{C}\in\mathbb{R}^{D\times D}$ the autocorrelation matrix, $\boldsymbol{E}=\operatorname{diag}(\boldsymbol{d})$, $\boldsymbol{d}\in\mathbb{R}_{+}^{DB}$ the diagonal elements. The bottleneck of evaluating this NLL lies on the calculation of the inverse and determinant of $\boldsymbol{\Sigma}$. Here, we can simplify the calculation using the Sherman–Morrison–Woodbury identity (matrix inversion lemma) and the companion matrix determinant lemma:
$\boldsymbol{\Sigma}^{-1}$ $=\left(\boldsymbol{E}+\boldsymbol{L}\left(\boldsymbol{C}\otimes\mathbf{I}_{R}\right)\boldsymbol{L}^{\top}\right)^{-1}$ (16)
$=\boldsymbol{E}^{-1}-\boldsymbol{E}^{-1}\boldsymbol{L}\left(\left(\boldsymbol{C}\otimes\mathbf{I}_{R}\right)^{-1}+\boldsymbol{L}^{\top}\boldsymbol{E}^{-1}\boldsymbol{L}\right)^{-1}\boldsymbol{L}^{\top}\boldsymbol{E}^{-1}.$
Consequently, the mahalanobis term in Eq. (15) becomes:
$\boldsymbol{\eta}^{\top}\boldsymbol{\Sigma}^{-1}\boldsymbol{\eta}=$ $\boldsymbol{\eta}^{\top}\boldsymbol{E}^{-1}\boldsymbol{\eta}-\boldsymbol{\eta}^{\top}\boldsymbol{E}^{-1}\boldsymbol{L}\left(\left(\boldsymbol{C}\otimes\mathbf{I}_{R}\right)^{-1}+\boldsymbol{L}^{\top}\boldsymbol{E}^{-1}\boldsymbol{L}\right)^{-1}\boldsymbol{L}^{\top}\boldsymbol{E}^{-1}\boldsymbol{\eta}$ (17)
$=$ $\boldsymbol{\eta}^{\top}\boldsymbol{E}^{-1}\boldsymbol{\eta}-\boldsymbol{\eta}^{\top}\boldsymbol{E}^{-1}\boldsymbol{L}\left(\boldsymbol{L}_{cap}\boldsymbol{L}_{cap}^{\top}\right)^{-1}\boldsymbol{L}^{\top}\boldsymbol{E}^{-1}\boldsymbol{\eta}$
$=$ $\boldsymbol{\eta}^{\top}\boldsymbol{E}^{-1}\boldsymbol{\eta}-\left(\boldsymbol{L}_{cap}^{-1}\boldsymbol{L}^{\top}\boldsymbol{E}^{-1}\boldsymbol{\eta}\right)^{\top}\left(\boldsymbol{L}_{cap}^{-1}\boldsymbol{L}^{\top}\boldsymbol{E}^{-1}\boldsymbol{\eta}\right)$
$=$ $\boldsymbol{\eta}^{\top}\boldsymbol{E}^{-1}\boldsymbol{\eta}-\boldsymbol{k}^{\top}\boldsymbol{k},$
where $\boldsymbol{k}=\boldsymbol{L}_{cap}^{-1}\boldsymbol{L}^{\top}\boldsymbol{E}^{-1}\boldsymbol{\eta}$, $\boldsymbol{L}_{cap}$ is the Cholesky factor of the capacitance matrix $\left(\left(\boldsymbol{C}\otimes\mathbf{I}_{R}\right)^{-1}+\boldsymbol{L}^{\top}\boldsymbol{E}^{-1}\boldsymbol{L}\right)$, the computing of $\boldsymbol{k}=\boldsymbol{L}_{cap}^{-1}\boldsymbol{L}^{\top}\boldsymbol{E}^{-1}\boldsymbol{\eta}$ can be fastly resolved by solving a linear system of equations $\boldsymbol{L}_{cap}\boldsymbol{k}=\boldsymbol{L}^{\top}\boldsymbol{E}^{-1}\boldsymbol{\eta}$. Since $\boldsymbol{E}$ is a diagonal matrix, the only matrix inverse we need to calculate in Eq. (17) is $\left(\boldsymbol{C}\otimes\mathbf{I}_{R}\right)^{-1}$, which can be further simplified as $\boldsymbol{C}^{-1}\otimes\mathbf{I}_{R}$. Recall $\boldsymbol{C}$ is a $D\times D$ autocorrelation matrix. Therefore, it is much easier than calculating the inverse of $\boldsymbol{\Sigma}$ which is a $DB\times DB$ matrix, and the computational cost does not scale with the number of time series $B$ in a batch.
The calculation of the determinant can also be greatly simplified with our parameterization:
$\ln\lvert\boldsymbol{\Sigma}\rvert$ $=\ln\lvert\boldsymbol{E}+\boldsymbol{L}\left(\boldsymbol{C}\otimes\mathbf{I}_{R}\right)\boldsymbol{L}^{\top}\rvert$ (18)
$=\ln\lvert\left(\boldsymbol{C}\otimes\mathbf{I}_{R}\right)^{-1}+\boldsymbol{L}^{\top}\boldsymbol{E}^{-1}\boldsymbol{L}\rvert+\ln\lvert\boldsymbol{C}\otimes\mathbf{I}_{R}\rvert+\ln\lvert\boldsymbol{E}\rvert$
$=2\sum^{DR}_{i}\ln\left[\boldsymbol{L}_{cap}\right]_{i,i}+2R\sum^{D}_{i}\ln\left[\boldsymbol{L}_{C}\right]_{i,i}+\sum^{DB}_{i}\ln\left[\boldsymbol{E}\right]_{i,i},$
where $\boldsymbol{L}_{C}$ is the Cholesky factor of the autocorrelation matrix $\boldsymbol{C}$.
A.3 Training Procedure
Compute used All models in the paper were trained in an Anaconda environment with access to one AMD Ryzen Threadripper PRO 5955WX CPU and four NVIDIA RTX A5000 GPU (24 GB of memory).
Batch size We adopt the approach of GPVar [2] by using $B=20$ time series in a sample slice and a batch size of 16. Given that our data sampler selects one slice of time series as a batch instead of sampling 16 slices simultaneously, we set $\mathtt{accumulate\_grad\_batches}$ to 16 to achieve an effective batch size of 16.
Training loop Each epoch of the training loop involves training the model with a maximum of 400 batches from the training set, followed by computing the NLL on the validation set. Training stops when any of these conditions are met:
  • We’ve reached a total of 10,000 gradient updates during model training,
  • We observed no improvement in the best value of the NLL on the validation set for 10 consecutive epochs.
We return the version of the model that achieved the best value of the NLL on the validation set.
A.4 Naive Baseline Description
In this paper, we employ VAR [33] (Vector Autoregression) and GARCH [34] (Generalized Autoregressive Conditionally Heteroscedastic) as two naive baseline models. The VAR($p$) model is defined as:
$\mathbf{z}_{t}=\mathbf{c}+\boldsymbol{A}_{1}\mathbf{z}_{t-1}+\dots+\boldsymbol{A}_{p}\mathbf{z}_{t-p}+\boldsymbol{\epsilon}_{t},\boldsymbol{\epsilon}_{t}\sim\mathcal{N}(\mathbf{0},\boldsymbol{\Sigma}_{\epsilon}),$ (19)
where $A_{i}$ is a $N\times N$ coefficient matrix and $\mathbf{c}$ is the intercept. We use a VAR model of lag 1 (i.e., a VAR(1) model) in the experiments. The parameters of Eq. (19) are estimated using ordinary least squares (OLS) following the procedure in [33].
The GARCH model describes the conditional covariance matrix of the current error term of a multivariate system. Suppose the model for the conditional mean is an AR($1$) model:
$\mathbf{z}_{t}=\mathbf{c}+\boldsymbol{A}_{1}\mathbf{z}_{t-1}+\boldsymbol{\epsilon}_{t},$ (20)
where the error term is modeled as:
$\boldsymbol{\epsilon}_{t}=\boldsymbol{H}_{t}^{1/2}\mathbf{e}_{t},$ (21)
where $\boldsymbol{H}_{t}$ is a $N\times N$ conditional covariance matrix and $\mathbf{e}_{t}$ is a $N\times 1$ standard normal vector $\mathbf{e}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{N})$. In the experiments, we use the DCC-GARCH(1,1)11(1,1)( 1 , 1 ) model [35], where the conditional covariance matrix $\boldsymbol{H}_{t}$ is defined as:
$\boldsymbol{H}_{t}=\boldsymbol{D}_{t}\boldsymbol{R}_{t}\boldsymbol{D}_{t},$ (22)
where $\boldsymbol{D}_{t}=\operatorname{diag}{(\mathbf{h}_{t})}^{1/2}$, and $\mathbf{h}_{t}$ contains the variances for each time series. $\boldsymbol{R}_{t}$ is the conditional correlation matrix in DCC-GARCH. The parameters of the DCC-GARCH model are estimated with the log-likelihood function:
$\mathcal{L}=-\frac{1}{2}\sum_{t=1}^{T}\left[N\ln{(2\pi)}+2\ln\lvert\boldsymbol{D}_{t}\rvert+\ln\lvert\boldsymbol{R}_{t}\rvert+\mathbf{e}_{t}^{\top}\boldsymbol{R}_{t}^{-1}\mathbf{e}_{t}\right].$ (23)
In this paper, we implement the VAR model with $\mathtt{statsmodels}$ [36] and the DCC-GARCH model with $\mathtt{mgarch}$ [37].
A.5 Hyperparameter Search
The hyperparameters and training configuration largely align with those in the GPVar paper [2]. The parameters of all DL models are learned using the Adam optimizer with $l2$ regularization set to 1e-8, and gradients are clipped to 10.0. Across all methods, we limit the total number of gradient updates to 10,000, and decay the learning rate by a factor of 2 after 500 consecutive updates without improvement. Table 3 lists the parameters that are tuned, along with the values of hyperparameters that are kept constant across all datasets and not subject to tuning.
Table 3: Hyperparameters values that are fixed or searched over a range during hyperparameter tuning.
Hyperparameter Value or Range Searched
learning rate [1e-4, 1e-4, 1e-2]
LSTM cells / $\mathtt{d\_model}$ of Transformer [10, 20, 40]
LSTM layers / Transformer decoder layers 2
$\mathtt{n\_heads}$ (Transformer) 2
rank 10
sampling dimension 20
dropout 0.01
batch size 16
To tune the hyperparameters of each model, we conduct a grid search of 9 parameters on each dataset. The best hyperparameters for a base model—dataset combination are selected based on the best validation loss. Once the optimal learning rate and hidden size are determined, we apply the same hyperparameters for models both with and without applying our approach.
The number of base kernels ($M$) for generating $\boldsymbol{C}_{t}$ and the associated lengthscale set $\{l_{m}\}_{m=1}^{M-1}$ are two additional hyperparameters when applying our method. The optimal values of $M$ and $\{l_{m}\}_{m=1}^{M-1}$ are selected in a similar way with hyperparameter search. The values of $M$ and $\{l_{m}\}_{m=1}^{M-1}$ searched during hyperparameter tuning are shown in Table 4. There are 6 possible combinations. For example, if we set $M=3$ and choose the lengthscale to start at $1.0$, the lengthscales for generating the component kernels will be $\{1.0,2.0\}$ since the last weight is for the identity matrix.
Table 4: Hyperparameters values of our method that are searched over a range during hyperparameter tuning.
Hyperparameter Value or Range Searched
number of kernels $M$ [2, 3, 4]
possible lengthscales $\{l_{m}\}_{m=1}^{M-1}$ [$\{0.5,1.5,\dots\}$, $\{1.0,2.0,\dots\}$]
A.6 Base Model Description and Input Features
The input to the base models consists of lagged time series values and generic features that encode time and time series identification. The number of lagged values used is determined by the time-frequency of each dataset. Specifically, we use [1, 24, 168] for hourly data, [1, 7, 14] for daily data, and [1, 2, 4, 12, 24, 48] for data with a granularity of less than one hour. For all other datasets, we only use the lag-1 values.
We use generic features to represent time. For datasets with a granularity of one hour or less, we include hour of the day and day of the week features. For daily datasets, we use the day of the week feature. Additionally, each time series is distinguished by an identifier number. All features are encoded with a single value; for example, the hour of the day feature takes values in [0, 23]. These feature values are concatenated with the RNN or Transformer input at each time step to generate the model input vector $\mathbf{y}_{t}$.
As illustrated in §4, our method requires a state vector $\mathbf{h}_{t}$ at each time step to generate the parameters for the predictive distribution and the dynamic weights for correlation matrix kernels. We use two different neural architectures for this purpose: RNN and Transformer, both of which preserve autoregressive properties. Specifically, we use LSTM as our base model for RNN and a decoder-only Transformer (i.e., the GPT model [22]) for the Transformer.
LSTM, a type of RNN architecture, is designed to model sequences and time series data. Unlike traditional RNNs, LSTMs can learn long-term dependencies, making them effective for tasks requiring context and memory over long sequences. A decoder-only Transformer is primarily used for sequence generation tasks, such as text generation, language modeling, and machine translation. It is a simplified version of the original Transformer model introduced by Vaswani et al. [38], including only the decoder part. The LSTM model can be formulated as:
$\mathbf{f}_{t}$ $=\sigma(\mathbf{W}_{f}\cdot[\mathbf{h}_{t-1},\mathbf{y}_{t}]+\mathbf{b}_{f}),$ (24)
$\mathbf{i}_{t}$ $=\sigma(\mathbf{W}_{i}\cdot[\mathbf{h}_{t-1},\mathbf{y}_{t}]+\mathbf{b}_{i}),$
$\tilde{\mathbf{C}}_{t}$ $=\tanh(\mathbf{W}_{C}\cdot[\mathbf{h}_{t-1},\mathbf{y}_{t}]+\mathbf{b}_{C}),$
$\mathbf{C}_{t}$ $=\mathbf{f}_{t}\odot\mathbf{C}_{t-1}+\mathbf{i}_{t}\odot\tilde{\mathbf{C}}_{t},$
$\mathbf{o}_{t}$ $=\sigma(\mathbf{W}_{o}\cdot[\mathbf{h}_{t-1},\mathbf{y}_{t}]+\mathbf{b}_{o}),$
$\mathbf{h}_{t}$ $=\mathbf{o}_{t}\odot\tanh(\mathbf{C}_{t}),$
where $\mathbf{y}_{t}$ is the input at each time step. The decoder-only Transformer can be formulated as:
$\mathbf{Q}$ $=\mathbf{Y}_{t}\mathbf{W}_{Q},$ (25)
$\mathbf{K}$ $=\mathbf{Y}_{t}\mathbf{W}_{K},$
$\mathbf{V}$ $=\mathbf{Y}_{t}\mathbf{W}_{V},$
$\mathbf{M}$ $=\text{Mask}(\mathbf{K}),$
$\mathbf{Z}$ $=\text{Softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{d_{k}}}+\mathbf{M}\right)\mathbf{V},$
$\mathbf{H}_{t}$ $=\text{LayerNorm}(\mathbf{Y}_{t}+\mathbf{Z}),$
$\mathbf{FFN}$ $=\text{ReLU}(\mathbf{H}_{t}\mathbf{W}_{1}+\mathbf{b}_{1})\mathbf{W}_{2}+\mathbf{b}_{2},$
$\mathbf{H}_{t}$ $=\text{LayerNorm}(\mathbf{H}_{t}+\mathbf{FFN}),$
where $\mathbf{H}_{t}$ is the output containing state vectors for all time steps, and $\mathbf{M}$ is a square causal mask for the sequence to preserve autoregressive properties.
Appendix B Metrics and Additional Results
B.1 Metric Definition
In this paper, we repeat the evaluation process on the testing set 10 times to compute the mean and standard deviation of all metrics. Metrics calculated in each independent evaluation process are based on the average results from all forecast instances in the testing set. For example, the $\operatorname{CRPS}_{\text{sum}}$ reported for $\mathtt{traffic}$ will be the average $\operatorname{CRPS}_{\text{sum}}$ of 7 forecast instances in its testing set. 100 samples of prediction were drawn for all evaluation processes.
B.1.1 Continuous Ranked Probability Score
The continuous ranked probability score ($\operatorname{CRPS}$) is defined as:
$\operatorname{CRPS}\left(F,z\right)=\operatorname{\mathbb{E}}_{F}\lvert Z-z\rvert-\frac{1}{2}\operatorname{\mathbb{E}}_{F}\lvert Z-Z^{\prime}\rvert,$ (26)
where $F$ is the cumulative distribution function (CDF) of the predicted variable, $z$ is the observation, $Z$ and $Z^{\prime}$ are independent copies of a set of prediction samples associated with the distribution $F$. For a single forecast instance, we calculate the average $\operatorname{CRPS}$ across time series and over the prediction horizon:
$\operatorname{\mathbb{E}}_{i,t}\left[\operatorname{CRPS}\left(F_{i,t},z_{i,t}\right)\right],$ (27)
where we use the empirical CDF to represent $F_{i,t}$ when predicting $z_{i,t}$. Since $\operatorname{CRPS}$ only compares a single ground-truth value to its predicted distribution, we also calculate the $\operatorname{CRPS}_{\text{sum}}$ [2, 16, 16] to assess multivariate dependencies in the time series data. $\operatorname{CRPS}_{\text{sum}}$ is computed by summing both the forecast and ground-truth values across all time series and then calculating the $\operatorname{CRPS}$ over the resulting sums:
$\operatorname{\mathbb{E}}_{t}\left[\operatorname{CRPS}\left(F_{t},\sum_{i}z_{i,t}\right)\right],$ (28)
where the empirical $F_{t}$ is obtained by summing samples across time series.
B.1.2 Quantile Loss
The quantile loss ($\rho$-risk) is another metric used in [9] to evaluate the performance of probabilistic forecasting:
$L_{\rho}\left(z,\hat{z}^{\rho}\right)=2\left(\hat{z}^{\rho}-z\right)\left(\left(1-\rho\right)\mathrm{I}_{\hat{z}^{\rho}>z}-\rho\mathrm{I}_{\hat{z}^{\rho}\leq z}\right),$ (29)
where $\mathrm{I}$ is a binary indicator function that equals 1 when the condition is met, $\hat{z}^{\rho}$ represents the predicted $\rho$-quantile, and $z$ represents the ground truth value. The quantile loss serves as a metric to assess the accuracy of a given quantile, denoted by $\rho$, from the predictive distribution. We summarize the quantile losses for the testing set across all time series segments by evaluating a normalized summation of these losses: $\left(\sum_{i,t}L_{\rho}\left(z_{i,t},\hat{z}_{i,t}^{\rho}\right)\right)/\left(\sum_{i,t}z_{i,t}\right)$. In this paper, we evaluate the $0.5$-risk and the $0.9$-risk following Salinas et al. [9].
B.1.3 Energy Score
The energy score ($\operatorname{ES}$) generalizes the $\operatorname{CRPS}$ to evaluate distributional forecasts of a vector-valued random variate, and thus another multivariate metric used in this paper:
$\operatorname{ES}(P,\mathbf{z})=\mathop{\operatorname{\mathbb{E}}}_{\boldsymbol{Z}\sim P}\lVert\boldsymbol{Z}-\mathbf{z}\rVert_{2}^{\beta}-\frac{1}{2}\mathop{\operatorname{\mathbb{E}}}_{\begin{subarray}{c}\boldsymbol{Z}\sim P\\ \boldsymbol{Z}^{\prime}\sim P\end{subarray}}\lVert\boldsymbol{Z}-\boldsymbol{Z}^{\prime}\rVert_{2}^{\beta},$ (30)
where $\lVert\mathbf{z}\rVert_{2}$ is the Euclidean norm and we use $\beta=1$ following [16] in this paper. Since we also want to aggregate over the prediction horizon, we calculate the Frobenius norm of the matrix $\lVert\mathbf{z}_{t+1:t+Q}\rVert_{F}$ in practice.
B.1.4 Root Relative Mean Squared Error
The root relative mean squared error ($\operatorname{RRMSE}$) is a metric commonly used for point forecast [39, 27, 40]. $\operatorname{RRMSE}$ is defined as:
$\operatorname{RRMSE}=\frac{\sqrt{\sum_{t=1}^{Q}\lVert\mathbf{z}_{t}-\hat{\mathbf{z}}_{t}\rVert_{2}^{2}}}{\sqrt{\sum_{t=1}^{Q}\lVert\mathbf{z}_{t}-\bar{\mathbf{z}}\rVert_{2}^{2}}},$ (31)
where $\hat{\mathbf{z}}_{t}$ is obtained by taking the mean of our prediction samples, $\bar{\mathbf{z}}$ is the mean value of the whole forecast instance. We use this metric to evaluate the mean prediction performance of our model.
B.2 Results on Other Forecasting Metrics
The results for $\operatorname{CRPS}$ (Table 5), $0.5$-risk (Table 6), $0.9$-risk (Table 7), $\operatorname{ES}$ (Table 8), and $\operatorname{RRMSE}$ (Table 9) are presented. An "N/A" entry in the tables indicates that the naive baseline models could not fit properly for this dataset. We observe consistent performance improvements in the base models using our method across different evaluation metrics. Notably, in the multivariate metric $\operatorname{ES}$, our method shows significant improvement, with an average increase of 5.58% for GPVar and 3.21% for Transformer.
Table 5: Comparison of $\operatorname{CRPS}$ accuracy. “w/o” denotes methods without time-dependent errors, while “w/” indicates our method. Boldface values indicate that models considering time-dependent errors have better performance. Mean and standard deviation are obtained from 10 runs of each model.
VAR GARCH GPVar Transformer
w/o w/ w/o w/
$\mathtt{exchange\_rate}$ 0.0070±0.0000 0.0438±0.0001 0.0171±0.0004 0.0141±0.0003 0.0092±0.0002 0.0081±0.0001
$\mathtt{solar}$ 0.9566±0.0022 0.9193±0.0010 0.7097±0.0047 0.7521±0.0027 0.5981±0.0021 0.5627±0.0018
$\mathtt{electricity}$ 0.1548±0.0003 0.2778±0.0010 0.0586±0.0004 0.0568±0.0002 0.0665±0.0003 0.0775±0.0001
$\mathtt{traffic}$ 19.9208±0.0495 0.4063±0.0002 0.1474±0.0001 0.1296±0.0001 0.1260±0.0001 0.1318±0.0001
$\mathtt{wiki}$ 334.6021±0.4936 3.0351±0.0048 0.3712±0.0003 0.3705±0.0004 0.3737±0.0003 0.2937±0.0002
$\mathtt{m4\_hourly}$ 0.2837±0.0004 0.3567±0.0004 0.1174±0.0002 0.1237±0.0002 0.1306±0.0002 0.1189±0.0002
$\mathtt{m1\_quarterly}$ N/A N/A 0.3942±0.0030 0.3538±0.0017 0.4448±0.0027 0.4367±0.002
$\mathtt{pems03}$ 0.1144±0.0001 0.3533±0.0002 0.0828±0.0000 0.0835±0.0001 0.0826±0.0001 0.0735±0.0000
$\mathtt{uber\_hourly}$ N/A N/A 0.1488±0.0003 0.1468±0.0002 0.1576±0.0003 0.1762±0.0003
avg. rel. impr. 3.59% avg. rel. impr. 3.13%
Table 6: Comparison of $0.5$-risk accuracy. “w/o” denotes methods without time-dependent errors, while “w/” indicates our method. Boldface values indicate that models considering time-dependent errors have better performance. Mean and standard deviation are obtained from 10 runs of each model.
VAR GARCH GPVar Transformer
w/o w/ w/o w/
$\mathtt{exchange\_rate}$ 0.0049±0.0000 0.0256±0.0001 0.0109±0.0003 0.0095±0.0004 0.0060±0.0001 0.0056±0.0001
$\mathtt{solar}$ 0.6140±0.0025 0.5621±0.0008 0.4998±0.0025 0.5246±0.0016 0.4233±0.0017 0.3958±0.0013
$\mathtt{electricity}$ 0.1113±0.0005 0.2014±0.0010 0.0405±0.0003 0.0397±0.0002 0.0449±0.0002 0.0505±0.0001
$\mathtt{traffic}$ 10.2654±0.0268 0.2722±0.0002 0.0933±0.0001 0.0859±0.0001 0.0803±0.0001 0.0794±0.0001
$\mathtt{wiki}$ 171.5009±0.2573 0.7225±0.0067 0.2231±0.0005 0.2236±0.0006 0.2030±0.0005 0.1487±0.0003
$\mathtt{m4\_hourly}$ 0.1992±0.0003 0.2365±0.0005 0.0807±0.0001 0.0849±0.0001 0.0880±0.0002 0.0808±0.0001
$\mathtt{m1\_quarterly}$ N/A N/A 0.2196±0.0023 0.1948±0.0005 0.2328±0.0008 0.2327±0.0014
$\mathtt{pems03}$ 0.0784±0.0001 0.2028±0.0002 0.0568±0.0000 0.0574±0.0001 0.0569±0.0000 0.0506±0.0000
$\mathtt{uber\_hourly}$ N/A N/A 0.1035±0.0002 0.1013±0.0002 0.1093±0.0003 0.1234±0.0002
avg. rel. impr. 2.75% avg. rel. impr. 3.88%
Table 7: Comparison of $0.9$-risk accuracy. “w/o” denotes methods without time-dependent errors, while “w/” indicates our method. Boldface values indicate that models considering time-dependent errors have better performance. Mean and standard deviation are obtained from 10 runs of each model.
VAR GARCH GPVar Transformer
w/o w/ w/o w/
$\mathtt{exchange\_rate}$ 0.0021±0.0000 0.0070±0.0000 0.0042±0.0001 0.0057±0.0001 0.0030±0.0000 0.0023±0.0001
$\mathtt{solar}$ 0.4676±0.0016 0.4393±0.0008 0.1617±0.0004 0.1597±0.0003 0.2744±0.0015 0.2710±0.0015
$\mathtt{electricity}$ 0.0414±0.0003 0.0744±0.0003 0.0211±0.0004 0.0185±0.0002 0.0281±0.0002 0.0366±0.0002
$\mathtt{traffic}$ 11.0170±0.0405 0.1689±0.0001 0.0666±0.0001 0.0580±0.0001 0.0609±0.0001 0.0698±0.0000
$\mathtt{wiki}$ 174.0756±0.3770 1.5906±0.0044 0.2136±0.0002 0.2048±0.0001 0.2117±0.0006 0.1764±0.0003
$\mathtt{m4\_hourly}$ 0.1029±0.0003 0.1309±0.0003 0.0452±0.0002 0.0463±0.0001 0.0525±0.0001 0.0475±0.0002
$\mathtt{m1\_quarterly}$ N/A N/A 0.3049±0.0044 0.2787±0.0027 0.3784±0.0031 0.3621±0.0037
$\mathtt{pems03}$ 0.0399±0.0000 0.1783±0.0001 0.0317±0.0000 0.0317±0.0001 0.0304±0.0000 0.0269±0.0000
$\mathtt{uber\_hourly}$ N/A N/A 0.0533±0.0002 0.0528±0.0001 0.0562±0.0001 0.0638±0.0002
avg. rel. impr. 0.22% avg. rel. impr. 0.91%
Table 8: Comparison of $\operatorname{ES}$ accuracy. “w/o” denotes methods without time-dependent errors, while “w/” indicates our method. Boldface values indicate that models considering time-dependent errors have better performance. Mean and standard deviation are obtained from 10 runs of each model.
VAR GARCH GPVar Transformer
w/o w/ w/o w/
$\mathtt{exchange\_rate}$ 0.1301±0.0002 0.6085±0.0009 0.3674±0.0067 0.2613±0.0047 0.1798±0.0039 0.1438±0.0026
$\mathtt{solar}$ (×$\times 10^{3}$) 1.7429±0.0043 1.7758±0.0015 1.6052±0.0095 1.6591±0.0050 1.5307±0.0049 1.4633±0.0044
$\mathtt{electricity}$ (×$\times 10^{5}$) 1.0102±0.0052 1.9422±0.0127 0.3569±0.0050 0.3172±0.0028 0.4031±0.0040 0.4754±0.0025
$\mathtt{traffic}$ 3.3585±0.010 (×$\times 10^{3}$) 4.4198±0.0020 2.4008±0.0020 2.2408±0.0015 2.2240±0.0021 2.2566±0.0014
$\mathtt{wiki}$ (×$\times 10^{7}$) 970.0242±2.5944 2.8857±0.0783 0.1149±0.0027 0.1155±0.0031 0.1236±0.004 0.1075±0.0046
$\mathtt{m4\_hourly}$ (×$\times 10^{3}$) 4.5109±0.0084 5.1849±0.0089 2.2729±0.0062 2.3611±0.0060 2.5877±0.0098 2.3440±0.0081
$\mathtt{m1\_quarterly}$ (×$\times 10^{2}$) N/A N/A 3.7565±0.0294 3.3676±0.0147 4.2149±0.0252 4.1596±0.0248
$\mathtt{pems03}$ (×$\times 10^{3}$) 1.3951±0.0009 5.4642±0.0067 1.0535±0.0010 1.0736±0.0015 1.0673±0.0012 0.9394±0.0004
$\mathtt{uber\_hourly}$ (×$\times 10^{3}$) N/A N/A 0.9035±0.0041 0.8773±0.0027 0.9377±0.0035 1.0566±0.0033
avg. rel. impr. 5.58% avg. rel. impr. 3.21%
Table 9: Comparison of $\operatorname{RRMSE}$ accuracy. “w/o” denotes methods without time-dependent errors, while “w/” indicates our method. Boldface values indicate that models considering time-dependent errors have better performance. Mean and standard deviation are obtained from 10 runs of each model.
VAR GARCH GPVar Transformer
w/o w/ w/o w/
$\mathtt{exchange\_rate}$ 0.0247±0.0000 0.0983±0.0002 0.0699±0.0012 0.0501±0.0010 0.0350±0.0008 0.0265±0.0007
$\mathtt{solar}$ 0.9365±0.0025 0.9556±0.0008 0.8195±0.0038 0.8334±0.0019 0.8114±0.0023 0.7761±0.0019
$\mathtt{electricity}$ 0.2732±0.0020 0.5584±0.0036 0.1010±0.0013 0.0912±0.0009 0.1130±0.0011 0.1293±0.0007
$\mathtt{traffic}$ 0.6312±0.0017 (×$\times 10^{3}$) 0.9894±0.0008 0.5383±0.0005 0.5061±0.0003 0.5025±0.0005 0.5052±0.0003
$\mathtt{wiki}$ 0.6519±0.0016 (×$\times 10^{4}$) 6.3386±0.2020 1.0288±0.0029 1.0393±0.0039 0.9292±0.0057 0.8752±0.0027
$\mathtt{m4\_hourly}$ 0.6163±0.0012 0.6848±0.0015 0.3072±0.0008 0.3168±0.0007 0.3420±0.0011 0.3179±0.0010
$\mathtt{m1\_quarterly}$ N/A N/A 19.1005±0.1246 17.0277±0.0845 20.2333±0.0830 20.2708±0.0924
$\mathtt{pems03}$ 0.3727±0.0003 0.8824±0.0013 0.2796±0.0003 0.2877±0.0005 0.2841±0.0003 0.2502±0.0001
$\mathtt{uber\_hourly}$ N/A N/A 0.2358±0.0012 0.2282±0.0008 0.2458±0.0010 0.2768±0.0009
avg. rel. impr. 5.48% avg. rel. impr. 2.85%
B.3 Training Dynamics
In Fig. 4 and Fig. 5, we compare the training dynamics of the base models trained with and without our method. Note that the likelihood losses of the two methods are not directly comparable, even for the same dataset and base model, due to differences in likelihood structure. We observe that while our method introduces more complexity into the likelihood loss, there is no evidence that it prolongs model convergence. On the contrary, our method can speed up convergence for some datasets in terms of the training steps used. We also report the training time in Table 10.
Table 10: Training cost comparison. “w/o” denotes methods without time-dependent errors, while “w/” indicates our method.
GPVar Transformer
w/o w/ w/o w/
sec./epoch epochs sec./epoch epochs sec./epoch epochs sec./epoch epochs
$\mathtt{exchange\_rate}$ 4.60 56 200.27 39 9.73 57 206.49 41
$\mathtt{solar}$ 6.18 39 74.16 51 15.37 121 181.26 132
$\mathtt{electricity}$ 7.44 71 119.06 94 19.38 63 103.15 65
$\mathtt{traffic}$ 10.50 55 225.20 48 28.30 84 247.58 100
$\mathtt{wiki}$ 12.45 30 164.81 33 28.16 51 351.06 48
$\mathtt{m4\_hourly}$ 7.42 67 189.14 43 17.81 43 355.27 65
$\mathtt{m1\_quarterly}$ 4.24 51 25.82 12 9.84 29 24.99 17
$\mathtt{pems03}$ 11.75 62 143.34 57 36.31 78 88.05 53
$\mathtt{uber\_hourly}$ 6.82 41 174.90 57 17.28 35 188.08 65
Figure 4: Training loss/validation loss vs training time of the GPVar model. “w/o” denotes methods without time-dependent errors, while “w/” indicates our method.
Figure 5: Training loss/validation loss vs training time of the Transformer model. “w/o” denotes methods without time-dependent errors, while “w/” indicates our method.
B.4 Additional Model Interpretation
In this section, we provide further insights into how our method improves the base model. We illustrate these insights by comparing the cross-correlation of the residuals from models with and without our method. Additionally, we demonstrate the performance of our method over the prediction horizon in multistep-ahead forecasting.
B.4.1 Comparison of Residual Correlation
Recall that our method models both the autocovariance of errors $\operatorname{Cov}(\eta_{i,t-\Delta},\eta_{i,t})$ and the cross-lag covariance $\operatorname{Cov}(\eta_{i,t-\Delta},\eta_{j,t})$ between all pairs of components in the multivariate series. With the calibration process introduced in §4.2, our method is expected to reduce error cross-correlation (autocorrelation and cross-lag correlation). Here, we compare the empirical ACF of the residuals $\eta_{i,t}$ of a single time series $i$, as well as the empirical cross-correlations of $\boldsymbol{\eta}_{t}$ across multiple time series.
We begin by comparing the ACF of the one-step-ahead prediction residuals with and without our method. The comparisons are provided for the following datasets: $\mathtt{solar}$ (Fig. 6), $\mathtt{electricity}$ (Fig. 7), $\mathtt{traffic}$ (Fig. 8), $\mathtt{wiki}$ (Fig. 9), $\mathtt{m4\_hourly}$ (Fig. 10), $\mathtt{pems03}$ (Fig. 11), and $\mathtt{uber\_hourly}$ (Fig. 12). We observe that the autocorrelation of the residuals is reduced after applying our method.
Next we compare cross-correlations of the one-step-ahead prediction residuals with and without our method. The comparisons are provided for the following datasets: $\mathtt{electricity}$ (Fig. 13), $\mathtt{traffic}$ (Fig. 14), $\mathtt{wiki}$ (Fig. 15), $\mathtt{m4\_hourly}$ (Fig. 16), $\mathtt{pems03}$ (Fig. 17), and $\mathtt{uber\_hourly}$ (Fig. 18). We also observe that the cross-correlations of the residuals are reduced after applying our method.
Figure 6: ACF comparison of the one-step-ahead prediction residuals with and without our method. The results depict the prediction outcomes generated by GPVar for four time series in the $\mathtt{solar}$ dataset.
Figure 7: ACF comparison of the one-step-ahead prediction residuals with and without our method. The results depict the prediction outcomes generated by GPVar for four time series in the $\mathtt{electricity}$ dataset.
Figure 8: ACF comparison of the one-step-ahead prediction residuals with and without our method. The results depict the prediction outcomes generated by GPVar for four time series in the $\mathtt{traffic}$ dataset.
Figure 9: ACF comparison of the one-step-ahead prediction residuals with and without our method. The results depict the prediction outcomes generated by GPVar for four time series in the $\mathtt{wiki}$ dataset.
Figure 10: ACF comparison of the one-step-ahead prediction residuals with and without our method. The results depict the prediction outcomes generated by GPVar for four time series in the $\mathtt{m4\_hourly}$ dataset.
Figure 11: ACF comparison of the one-step-ahead prediction residuals with and without our method. The results depict the prediction outcomes generated by GPVar for four time series in the $\mathtt{pems03}$ dataset.
Figure 12: ACF comparison of the one-step-ahead prediction residuals with and without our method. The results depict the prediction outcomes generated by GPVar for four time series in the $\mathtt{uber\_hourly}$ dataset.
Figure 13: Cross-correlation comparison of the one-step-ahead prediction residuals with and without our method. The results depict the prediction outcomes generated by GPVar for four time series in the $\mathtt{electricity}$ dataset.
Figure 14: Cross-correlation comparison of the one-step-ahead prediction residuals with and without our method. The results depict the prediction outcomes generated by GPVar for four time series in the $\mathtt{traffic}$ dataset.
Figure 15: Cross-correlation comparison of the one-step-ahead prediction residuals with and without our method. The results depict the prediction outcomes generated by GPVar for four time series in the $\mathtt{wiki}$ dataset.
Figure 16: Cross-correlation comparison of the one-step-ahead prediction residuals with and without our method. The results depict the prediction outcomes generated by GPVar for four time series in the $\mathtt{m4\_hourly}$ dataset.
Figure 17: Cross-correlation comparison of the one-step-ahead prediction residuals with and without our method. The results depict the prediction outcomes generated by GPVar for four time series in the $\mathtt{pems03}$ dataset.
Figure 18: Cross-correlation comparison of the one-step-ahead prediction residuals with and without our method. The results depict the prediction outcomes generated by GPVar for four time series in the $\mathtt{uber\_hourly}$ dataset.
B.4.2 Performance Breakdown at Each Forecast Step
To investigate our performance gain at each forecast step, we calculate the $\operatorname{CRPS}_{\text{sum}}$ for each forecast step. The results are shown in Fig. 19 (GPVar) and Fig. 20 (Transformer). Note that the $\operatorname{CRPS}_{\text{sum}}$ reported in this section may have different scales compared to previous sections since they are not normalized. Since the predicted values are used as inputs within the prediction range during multistep-ahead forecasting, the residuals accumulate the effects of inaccuracies from previous steps. Therefore, the performance improvement depends not only on our modeling of error correlations but also on the properties of the residuals. These properties can be influenced by the absolute and relative time of the forecast and the seasonality of the data. For data without strong seasonality, residuals tend to be larger when predicting further ahead, making error accumulation more apparent. In contrast, for data with strong seasonality, the impact can vary. We can observe that, in most scenarios, $\operatorname{CRPS}_{\text{sum}}$ is reduced at the early forecasting stages. As predictions extend further into the future, some datasets (e.g., $\mathtt{traffic}$ in Fig. 19) show decreased improvement, likely due to seasonality effects. In contrast, other datasets (e.g., $\mathtt{wiki}$ in Fig. 19) exhibit larger improvements further into the future, possibly because the residuals accumulate over the steps.
Figure 19: Step-wise $\operatorname{CRPS}_{\text{sum}}$ accuracy of GPVar. “w/o” denotes methods without time-dependent errors, while “w/” indicates our method.
Figure 20: Step-wise $\operatorname{CRPS}_{\text{sum}}$ accuracy of Transformer. “w/o” denotes methods without time-dependent errors, while “w/” indicates our method.
B.5 Alternative Parametrization of $\boldsymbol{C}_{t}$
B.5.1 Learnable Lengthscales
In this paper, the lengthscales are fixed when generating the correlation matrix $\boldsymbol{C}_{t}$, and the flexibility of $\boldsymbol{C}_{t}$ comes from dynamically generating the component weights of the kernel matrices. Making these lengthscales learnable parameters to find the optimal set of $\{l_{m}\}_{m=1}^{M-1}$ is another approach we can explore to increase modeling flexibility. Based on the best model identified in Table 1, we experiment with making the lengthscales learnable parameters, jointly optimized with the base model. The results are shown in Table 11. We do not observe significant enhancement from learnable lengthscales.
Table 11: Comparison of $\operatorname{CRPS}_{\text{sum}}$ accuracy. “w/o” denotes methods without time-dependent errors, while “w/” indicates our method. “w/ (l)” indicates the lengthscales are learnable parameters. Boldface values indicate that models considering time-dependent errors have better performance. Mean and standard deviation are obtained from 10 runs of each model.
GPVar Transformer
w/o w/ w/ (l) w/o w/ w/ (l)
$\mathtt{exchange\_rate}$ 0.0068±0.0004 0.0117±0.0004 0.0045±0.0001 0.0055±0.0002 0.0042±0.0002 0.0072±0.0002
$\mathtt{solar}$ 0.7103±0.0065 0.6929±0.0039 0.7727±0.0040 0.4960±0.0034 0.4132±0.0027 0.4138±0.0023
$\mathtt{electricity}$ 0.0430±0.0005 0.0403±0.0004 0.0351±0.0003 0.0494±0.0004 0.0638±0.0003 0.0858±0.0006
$\mathtt{traffic}$ 0.1095±0.0002 0.0649±0.0002 0.1297±0.0003 0.0717±0.0002 0.0981±0.0002 0.0950±0.0002
$\mathtt{wiki}$ 0.1745±0.0008 0.0743±0.0009 0.4839±0.0021 0.0841±0.0013 0.0500±0.0005 0.0472±0.0004
$\mathtt{m4\_hourly}$ 0.0613±0.0004 0.0358±0.0002 0.0527±0.0003 0.0651±0.0004 0.0616±0.0003 0.0355±0.0003
$\mathtt{m1\_quarterly}$ 0.3942±0.0030 0.3538±0.0017 0.3534±0.0017 0.4448±0.0027 0.4367±0.0028 0.3709±0.0120
$\mathtt{pems03}$ 0.0503±0.0001 0.0491±0.0002 0.0456±0.0001 0.0490±0.0001 0.0386±0.0001 0.0330±0.0001
$\mathtt{uber\_hourly}$ 0.0342±0.0006 0.0222±0.0004 0.0218±0.0003 0.0632±0.0003 0.0513±0.0005 0.0969±0.0005
B.5.2 Using Autocorrelations of an AR$(p)$ process
One could parameterize $\boldsymbol{C}_{t}$ as fully-learnable positive definite symmetric Toeplitz matrices. For instance, an AR$(p)$ process has a autocorrelation matrix with a Toeplitz structure, allowing the modeling of negative correlations. This alternative approach may offer more flexibility in capturing complex correlation patterns in multivariate time series data. The autocorrelations of an AR$(p)$ process can be obtained by solving a set of equations known as Yule-Walker equations [41]. For example, if we consider an AR$(2)$ process and $\rho_{k}$ as the autocorrelation at lag $k$:
$z_{t}=\phi_{1}z_{t-1}+\phi_{2}z_{t-2}+\epsilon_{t}.$ (32)
where $\phi_{1}$ and $\phi_{2}$ are the coefficients. We have $\rho_{0}=1$ by definition and:
$\rho_{1}$ $=\phi_{1}\rho_{0}+\phi_{2}\rho_{1},$ (33)
$\ldots$
$\rho_{k}$ $=\phi_{1}\rho_{k-1}+\phi_{2}\rho_{k-2},k\geq 2.$
Since $\rho_{0}=1$, we can solve for $\rho_{1}$:
$\rho_{1}=\frac{\phi_{1}}{1-\phi_{2}},$ (34)
and for any $k\geq 2$, we can solve $\rho_{k}$ iteratively by:
$\rho_{k}=\phi_{1}\rho_{k-1}+\phi_{2}\rho_{k-2},k\geq 2.$ (35)
The collection $\{\rho_{0},\rho_{1},\dots,\rho_{k},\dots,\rho_{D-1}\}$ forms a row or column of a Toeplitz matrix and can be used to parameterize $\boldsymbol{C}_{t}$. We perform a hyperparameter search to find the best AR order $p$ based on the validation loss. As shown in Table 12, although the correlation matrix $\boldsymbol{C}_{t}$ parameterized by an AR process shows promise in modeling both positive and negative correlations, it does not empirically provide an overall enhancement compared to the kernel method used in this paper. This may be because cross-correlations in time series are predominantly positive. However, the AR method does show significant improvements on certain datasets where the kernel method does not perform well. For example, the AR method greatly improves GPVar on $\mathtt{exchange\_rate}$ and Transformer on $\mathtt{electricity}$.
Table 12: Comparison of $\operatorname{CRPS}_{\text{sum}}$ accuracy. “w/o” denotes methods without time-dependent errors, while “w/” indicates our method. “w/ (AR)” indicates $\boldsymbol{C}_{t}$ is parameterized by an AR process. Boldface values indicate that models considering time-dependent errors have better performance. Mean and standard deviation are obtained from 10 runs of each model.
GPVar Transformer
w/o w/ w/ (AR) w/o w/ w/ (AR)
$\mathtt{exchange\_rate}$ 0.0068±0.0004 0.0117±0.0004 0.0051±0.0002 0.0055±0.0002 0.0042±0.0002 0.0088±0.0004
$\mathtt{solar}$ 0.7103±0.0065 0.6929±0.0039 0.5923±0.0042 0.4960±0.0034 0.4132±0.0027 0.3362±0.0025
$\mathtt{electricity}$ 0.0430±0.0005 0.0403±0.0004 0.0433±0.0007 0.0494±0.0004 0.0638±0.0003 0.0252±0.0002
$\mathtt{traffic}$ 0.1095±0.0002 0.0649±0.0002 0.1095±0.0004 0.0717±0.0002 0.0981±0.0002 0.0878±0.0003
$\mathtt{wiki}$ 0.1745±0.0008 0.0743±0.0009 0.2375±0.0013 0.0841±0.0013 0.0500±0.0005 0.0512±0.0008
$\mathtt{m4\_hourly}$ 0.0613±0.0004 0.0358±0.0002 0.0298±0.0002 0.0651±0.0004 0.0616±0.0003 0.0680±0.0003
$\mathtt{m1\_quarterly}$ 0.3942±0.0030 0.3538±0.0017 0.1692±0.0029 0.4448±0.0027 0.4367±0.0028 0.4348±0.0028
$\mathtt{pems03}$ 0.0503±0.0001 0.0491±0.0002 0.0787±0.0002 0.0490±0.0001 0.0386±0.0001 0.0656±0.0001
$\mathtt{uber\_hourly}$ 0.0342±0.0006 0.0222±0.0004 0.0375±0.0004 0.0632±0.0003 0.0513±0.0005 0.0770±0.0007
B.6 Qualitative Results on Forecasting
In this section, we provide qualitative analysis of the actual prediction performance by visualizing the predictions.
Figure 21: Visualization of forecasting results on exchange_rate using GPVar with our method.
Figure 22: Visualization of forecasting results on $\mathtt{solar}$ using GPVar with our method.
Figure 23: Visualization of forecasting results on $\mathtt{electricity}$ using GPVar with our method.
Figure 24: Visualization of forecasting results on $\mathtt{traffic}$ using GPVar with our method.
Figure 25: Visualization of forecasting results on $\mathtt{wiki}$ using GPVar with our method.
Figure 26: Visualization of forecasting results on $\mathtt{m4\_hourly}$ using GPVar with our method.
Figure 27: Visualization of forecasting results on $\mathtt{pems03}$ using GPVar with our method.
Figure 28: Visualization of forecasting results on $\mathtt{uber\_hourly}$ using GPVar with our method.